Measurements are made of the intensity distribution within the central bright fringe in a Young's interference pattern. At a particular value of y, it is found that I/Imax= 0.810 when 600-nm light is used. What wavelength of light should be used to reduce the relative intensity at the same location to 64.0% of the maximum untensity?
1
Expert's answer
2020-06-15T10:27:35-0400

As per the question,

"\frac{I}{I_{max}}=0.810"

"\lambda_1=600nm =600\times 10^{-9}m"

"\lambda_2=?"

relative intensity at the same location "=64.0\%"

"I=I_{max}\cos^2({\frac{\pi yd}{\lambda_1 L})}"


"\Rightarrow \frac{\pi yd}{ L}=\lambda_1\cos^{-1}(\frac{I}{I_{max}})^{1/2}"


"\Rightarrow \frac{\pi y d}{L}=600 \cos^{-1}(0.810)^{1/2}"

"\Rightarrow \frac{\pi y d}{L}=271 nm"

Now,

"\frac{I}{I_{max}}=0.64"


"\lambda_2=\frac{\frac{\pi y d}{L}}{\cos^{-1}(\frac{I}{I_{max}})^{1/2}}=38.14nm"